1.Suppose that the width of a rectangle is 5 inches shorter than the length and that the perimeter of the rectangle is 50.
a) Set up an equation for the perimeter involving only L, the length of the rectangle.
Answer:
b) Solve this equation algebraically to find the length of the rectangle. Find the width as well.
2.5x+4y=12 for y
Answer:

Show your work here:
b) When graphed, this equation would be a line. By examining your answer to part a, what is the slope and y-intercept of this line?
Slope = ______
Y-intercept = _____
c) Using your answer from part a, find the corresponding value of y when x = 16.



Answer by stanbon(57387)

1.Suppose that the width of a rectangle is 5 inches shorter than the length and that the perimeter of the rectangle is 50.
a) Set up an equation for the perimeter involving only L, the length of the rectangle.
Let length be L ; then width = L-5
Perimeter = 2(length + width)
Answer: P(L) = 2(L + L-5) = 2(2L-5) = 4L-10 inches
--------------------------------------------------------------
b) Solve this equation algebraically to find the length of the rectangle. Find the width as well.
4L-10 = P
L = (P+10)/4
----------------------------------------

2.5x+4y=12 for y
4y = -5x + 12
Answer: y = (-5/4)x + 3
========================================
Show your work here:
b) When graphed, this equation would be a line. By examining your answer to part a, what is the slope and y-intercept of this line?
Slope = -5/4______
Y-intercept = 3
----_____
c) Using your answer from part a, find the corresponding value of y when x = 16.
y = (-5/4)*16 + 3
y = -20 + 3
y = -17
=================================
Cheers,
Stan H.